# Dense Feature Interaction Network for Image Inpainting Localization

### v1

@@ Abstract
Image inpainting, which is the task of filling in missing areas in an image, is a common image editing technique. Inpainting can be used to conceal or alter image contents in malicious manipulation of images, driving the need for research in image inpainting detection. Existing methods mostly rely on a basic encoder-decoder structure, which often results in a high number of false positives or misses the inpainted regions, especially when dealing with targets of varying semantics and scales. Additionally, the absence of an effective approach to capture boundary artifacts leads to less accurate edge localization. In this paper, we describe a new method for inpainting detection based on a Dense Feature Interaction Network (DeFI-Net). DeFI-Net uses a novel feature pyramid architecture to capture and amplify multi-scale representations across various stages, thereby improving the detection of image inpainting by better revealing feature-level interactions. Additionally, the network can adaptively direct the lower-level features, which carry edge and shape information, to refine the localization of manipulated regions while integrating the higher-level semantic features. Using DeFI-Net, we develop a method combining complementary representations to accurately identify inpainted areas. Evaluation on five image inpainting datasets demonstrate the effectiveness of our approach, which achieves state-of-the-art performance in detecting inpainting across diverse models.
@@ I Introduction
Image inpainting is a common image editing technique that restores damaged or missing regions with visually credible structures and textures based on the other part of the image. Inpainting has applications such as photo restoration, obstacle removal, and augmented reality. Traditional inpainting methods formulate it as an optimization problem [1], where the goal is to find the most suitable information to fill in the gaps while maintaining overall coherence in the reconstructed image. It is mainly divided into two branches: pixel propagation based [2, 3, 4], and image patches matching based [5, 6, 7, 8, 9]. The rapid advancements of learning-based methods, especially the recent generative AI tools such as OpenAI DALL-E [^1] https://openai.com/dall-e-2, Stable Diffusion [^2] https://huggingface.co/runwayml/stable-diffusion-inpainting, and Adobe Firefly [^3] https://www.adobe.com/products/firefly.html have enabled users to create and perform image editing, including inpainting, by simply providing some text prompts. The generative AI backed inpainting method exacerbates the negative impacts of malicious use of image inpainting as a means of disinformation. A notable recent example is the case where nudity images created with AI-powered image inpainting were spread for cyber targeting female students in a California middle school [^4] https://www.latimes.com/california/story/2024-02-26/beverly-hills-middle-school-is-the-latest-to-be-rocked-by-deepfake-scandal.
@@
Figure 1: Visualization of features at varying levels in describing inpainting clues. Upper feature maps illustrate a baseline network PSCC-Net [10] that struggles to capture discriminative patterns despite employing multi-level feature learning. In contrast, the lower maps show our DeFI-Net, leveraging dense feature interaction to reveal diverse and complementary indicators crucial for accurate inpainting detection.
@@
To date, many methods have been developed to expose local image manipulation [11, 12, 10, 13, 14, 15]. Yet locating inpainted regions is still a challenging problem, which can be attributed to the following two factors. First, unlike traditional copy-move and splicing manipulations, which often result in noticeable artifacts such as duplicated pixels or inconsistent patterns in semantic and color, advanced image inpainting algorithms mitigate signal and visual artifacts by preserving statistical properties [16] extracted not only from surrounding pixels but also inferred from massive datasets in a data-driven manner. Additionally, the varied sizes of inpainted regions, ranging from large object editing to small watermark, signature, or trademark removal, pose further challenges in the detection process.
@@
Although several studies have dedicated exclusively to identifying image inpainting, such as using high-pass fully convolutional network [17], Neural Architecture Search (NAS) based model[18], Transformer architecture[19], or primary–secondary network [20], most of them adopt conventional encoder-decoder architectures, prioritizing deep high-level555 In this paper, we use “level” to indicate the abstraction level of extracted information, “stage” to describe a group of layers in CNN models, and “layer” as the basic building blocks of CNN architecture. information while overlooking shallow features that contain crucial contextual details essential for uncovering inpainting artifacts, such as subtle inconsistencies in edge and shapes. Furthermore, for learning inconsistencies at inpainting boundaries, relying solely on ordinary convolution combinations and fixed edge extraction operators makes it challenging to capture representative artifact features. Such limitations reduce the model’s ability to address complex or small inpainting areas, resulting in false positives or diminished localization accuracy.
@@
Figure 2: A simplified schematic comparison between existing multi-scale learning strategies and our DeFI-Net. Unlike previous methods that rely on progressive interaction (e.g., PSCC-Net [10]), focus more on high-level feature fusion (e.g., MVSS-Net [14]), or low-level features (e.g., GCA-Net [21]), our model offers dense feature interaction and adaptively utilizes complementary features from all levels to enhance inpainting localization.
@@
We use the insight that features at various layers [22] within deep learning networks can expose the subtle inconsistencies introduced by the inpainting operation in a different while complementary manner (refer to Fig. 1). Low-level features are closer to the input layer and have smaller receptive fields, thus primarily capturing signal artifacts, like edges and corners lacking semantic information [23]. Features at the middle level represent more abstract patterns and object parts, while high-level semantic features describing entire objects, scenes, or concepts are derived from the network’s deeper layers. Our goal is to fully leverage and enhance the diverse contributions of these features towards localizing image inpainting. To this end, we develop a new method for inpainting detection based on a Dense Feature Interaction Network (DeFI-Net). DeFI-Net uses a feature pyramid architecture to capture and amplify multi-scale representations across various stages, which can more effectively reveal the feature-level interactions for detecting image inpainting. This structure adaptively directs the lower-level features, which carry edge and shape information, to refine the localization of manipulated edges while integrating the higher-level semantic features. Using DeFI-Net, we have developed a priori feature-guided edge attention module, incorporating inpainting edge prediction as an auxiliary task to enhance feature representation and improve localization accuracy at edge details. Evaluation on five image inpainting datasets demonstrate the effectiveness of our approach, which achieves state-of-the-art performance in detecting inpainting across diverse models. The main contributions of our work can be summarized as follows.
@@
    We introduce the DeFI-Net to highlight how layered features aid inpainting detection, with a dense feature pyramid architecture hierarchically and thoroughly encoding multi-view features through dense connections.
@@
    We further design a complementary supervision approach by adding edge supervision with reverse attention on lower-layer features to the inpainting region localization using multi-level features.
@@
    To effectively integrate multi-level features, we design a spatial weight learning module that adaptively assigns fusion weights to different levels of features, providing representative and compact contexts for recognizing inpainting.
@@ II-A Image inpainting manipulation
Figure 3: Architecture of the proposed DeFI-Net. Left: The Dense Feature Pyramid Learning module densely integrates multi-stage features extracted from the backbone network. Right: Structure of REAE and SA-FF modules for inpainting artifact enhancement and multi-level semantic fusion.
@@ II-B Image inpainting detection
Detecting of image inpainting is often solved with machine learning based on convolutional neural networks (CNNs) [41]. These CNN-based approaches exploit hierarchical neural networks to automatically learn discriminative features for detecting inpainted regions. Zhu et al. [42] pioneered CNN-based image inpainting detection by employing a label matrix to guide automatic feature learning and the weighted cross-entropy as the loss function. Several studies [43, 44, 45] employed Long Short-Term Memory (LSTM) models to learn artifacts for manipulation localization. HP-FCN [17] was proposed to enhance inpainting traces via high-pass filtering and learn discriminative features from image residuals. Similarly, IID-Net [18] used various combined filters to enhance inpainting features, and introduced a Neural Architecture Search (NAS) algorithm to adjust feature extraction cells. These conventional encoder-decoder structures without pooling layers may hinder the model’s capability to detect intricate tampering and result in spatial information loss. Several methods [46, 11, 47, 14] introduces multi-branch models to combine features from noise/frequency domains or patches [48]. New feature extraction models, including Transformer-based architectures [19, 49] or DenseNet-based approaches [50, 20] have been used for inpainting detection. Furthermore, researchers have explored multi-scale feature learning techniques. For example, MVSS-Net [13, 14] employed features learned from different stages of the backbone network and incorporated a two-branch supervision method considering RGB domain and noise patterns. Liu et al. [10] introduced PSCC-Net, leveraging the multi-stage features learned from the HRNet [51] for progressive mask prediction. GCA-Net [21] designed a gated context attention network to capture fine image discrepancies in inpainted images. To improve global context representations and minimize feature loss, they refer to the UNet++ [52] architecture, which utilizes multi-scale bottom-up feature propagation in the decoding process. Recently, Yu et al. introduced DiffForensics [15], which employs a self-supervised denoising diffusion paradigm integrated with an encoder-decoder architecture to detect and localize tampering.
@@ III DeFI-Net
In this work, we introduce the DeFI-Net for image inpainting localization, as shown in Fig. 3. Taking an RGB image $I\in\mathbb{R}^{H\times W\times C}$ as input, the model will output a single-channel edge mask $M_{e}\in\mathbb{R}^{H\times W\times 1}$ and region mask $M_{r}\in\mathbb{R}^{H\times W\times 1}$. The model begins with multiple stages of feature extraction from the backbone network. The multi-stage features are interacted and fused by three components: 1) a Dense Feature Pyramid Learning structure (DFPL) to introduce multi-scale features containing detailed texture, contextual and semantic information (see Sec. III-A); 2) a Reverse Edge-Attention Enhancement (REAE), with low-mid-level features as a prior guiding the model to focus on edge inconsistencies caused by inpainting (see Sec. III-B); 3) a Spatial Adaptive Feature Fusion (SA-FF) aims to learn the fusion weights of the high-level and mid-low-level features adaptively, unifying the representation of inpainting to achieve accurate detection results ( see Sec. III-C).
@@ III-A Dense Feature Pyramid Learning
Different layers in a Convolutional Neural Network (CNN) model have receptive fields of different sizes and extract image features of different levels of detail. Generally, early or shadow layers mostly extract low-level and generic image cues (e.g., edges and corners), and the middle layers carry shapes and textural patterns. In contrast, deeper layers capture high-level semantics or specific image cues [55, 44, 23]. Artifacts of image inpainting may be reflected at all three levels of features, each contributing in a complementary manner. For example, pixel filling can cause distortions or blur at edges, detectable by low-level features. Middle-level features mainly capture texture inconsistencies in the filled areas, like structural mismatches or color discordance. Differently, high-level features identify changes in an image’s semantic content, such as missing object parts or altered scene interpretation, aiding the model in identifying inpainting areas.
Our goal is to design an effective extraction and fusion of these complementary features (as shown in Fig. 1) to identify image inpainting. We first use the backbone network to extract multi-stage features from the input image. Using a CNN (e.g.,  [56]) divided into five stages as an example, we define the shallow features from Stage 1 and Stage 2 as low-level features $f_{1}$ and $f_{2}$, the features extracted from Stage 3 as mid-level feature $f_{3}$, while the more complex deep features from Stage 4 as high-level features $f_{4}$. Inspired by the Feature Pyramid Network (FPN) [22] and Path Aggregation Network (PANet) [57], we design a multi-directional, multi-scale fusion method, as shown in Fig. 4. Differently, we treat multi-level features as independent entities, emphasizing the specificity of features at different stages. Our densely connected fusion approach improves the efficiency of information interaction between adjacent and cross layers, progressively enhancing context awareness.
Taking a single-scale feature as an example, our DFPL module considers the features of the same scale and all the neighboring features to fully exploit and enhance the useful context information for image inpainting localization. Specifically, to obtain the feature $f_{i}^{j}$ (where $i\in[1,4]$ denotes the stage of the backbone to which the current feature belongs, and $j\in[1,3]$ represents the state of the feature in DFPL, $j$ as 1 for the initial state involving channel reduction, 2 for the bottom-up path, and 3 for the top-down path), we concatenate its neighboring features in the channel dimension and use $1\times 1$ convolution to integrate and select features. This process can be expressed by the equation below, where we have removed the max pooling, upsampling operations and cases with $i$ in the boundary condition for simplicity.
Figure 4: Comparison of our DFPL with existing Feature Pyramid Network (FPN) [22] in multi-level feature interaction.

### v2

@@ Abstract
Image inpainting, the process of filling in missing areas in an image, is a common image editing technique. Inpainting can be used to conceal or alter image contents in malicious manipulation of images, driving the need for research in image inpainting detection. Most existing methods use a basic encoder-decoder structure, which often results in a high number of false positives or misses the inpainted regions, especially when dealing with targets of varying semantics and scales. Additionally, the lack of an effective approach to capture boundary artifacts leads to less accurate edge localization. In this paper, we describe a new method for inpainting detection based on a Dense Feature Interaction Network (DeFI-Net). DeFI-Net uses a novel feature pyramid architecture to capture and amplify multi-scale representations across various stages, thereby improving the detection of image inpainting by better strengthening feature-level interactions. Additionally, the network can adaptively direct the lower-level features, which carry edge and shape information, to refine the localization of manipulated regions while integrating the higher-level semantic features. Using DeFI-Net, we develop a method combining complementary representations to accurately identify inpainted areas. Evaluation on seven image inpainting datasets demonstrates the effectiveness of our approach, which achieves state-of-the-art performance in detecting inpainting across diverse models. Code and models are available at https://github.com/Boombb/DeFI-Net_Inpainting.
@@ I Introduction
Image inpainting is a common image editing technique that restores damaged or missing regions with visually credible structures and textures from the surrounding image content. Inpainting has applications such as photo restoration, obstacle removal, and augmented reality. Traditional inpainting methods formulate it as an optimization problem [1], where the goal is to find the most suitable information to fill in the gaps while maintaining overall coherence in the reconstructed image. It is mainly divided into two branches: pixel propagation based [2, 3, 4], and image patches matching based [5, 6, 7, 8, 9]. The rapid advancements of learning-based methods, especially the recent generative AI tools such as OpenAI DALL-E [^1] https://openai.com/dall-e-2, Stable Diffusion [^2] https://huggingface.co/runwayml/stable-diffusion-inpainting, and Adobe Firefly [^3] https://www.adobe.com/products/firefly.html have enabled users to create and perform image editing, including inpainting, by simply providing some text prompts. The generative AI backed inpainting method exacerbates the negative impacts of malicious use of image inpainting as a means of disinformation. A notable recent example is the case where nude images created with AI-powered image inpainting were spread for cyber targeting female students in a California middle school [^4] https://www.latimes.com/california/story/2024-02-26/beverly-hills-middle-school-is-the-latest-to-be-rocked-by-deepfake-scandal.
@@
Figure 1: Visualization of features at varying levels in describing inpainting clues. Yellow regions indicate higher responsiveness, highlighting key regions for identifying inpainting, while blue regions represent low responsiveness. Upper feature maps illustrate a baseline network PSCC-Net [10] that struggles to capture discriminative patterns despite employing multi-level feature learning. In contrast, the lower maps show our DeFI-Net, leveraging dense feature interaction to reveal diverse and complementary indicators crucial for accurate inpainting detection.
@@
To date, many methods have been developed to expose local image manipulation [11, 12, 10, 13, 14, 15]. Yet locating inpainted regions is still a challenging problem, which can be attributed to the following two factors. First, unlike traditional copy-move and splicing manipulations, which often produce noticeable artifacts such as duplicated pixels or inconsistent patterns in semantic and color, advanced image inpainting algorithms mitigate signal and visual artifacts by preserving statistical properties [16] from both surrounding pixels and large-scale datasets in a data-driven manner. Additionally, the varied sizes of inpainted regions, ranging from large object editing to small watermark, signature, or trademark removal, pose further challenges in the detection process.
@@
Although several studies have been dedicated exclusively to identifying image inpainting, such as using high-pass fully convolutional network [17], Neural Architecture Search (NAS) based model [18], Transformer architecture [19], or primary–secondary network [20], most adopt conventional encoder-decoder architectures. These methods prioritize high-level555 In this paper, we use “level” to indicate the abstraction level of extracted information, “stage” to describe a resolution convolution stream, as defined in HRNet [21], and “layer” as the basic building blocks of CNN architecture. abstract information while overlooking the features that contain crucial contextual details, such as subtle inconsistencies in edges and shapes, which are essential for uncovering inpainting artifacts. Furthermore, for learning inconsistencies at inpainting boundaries, relying solely on ordinary convolution combinations and fixed edge extraction operators makes it challenging to capture representative artifact features. Such limitations reduce the model’s ability to address complex or small inpainting areas, resulting in false positives or diminished localization accuracy.
@@
Figure 2: A simplified schematic comparison between existing multi-scale learning strategies and our DeFI-Net. Unlike previous methods that rely on progressive interaction (e.g., PSCC-Net [10]), focus more on high-level feature fusion (e.g., MVSS-Net [14]), or low-level features (e.g., GCA-Net [22]), our model offers dense feature interaction and adaptively utilizes complementary features from all levels to enhance inpainting localization. The yellow, orange, and red circles represent low-, mid-, and high-level features, respectively. The black arrows indicate the flow of features through the model. Finally, depending on the model design, the output is directed to different objective functions.
@@
We use the insight that features at various layers [23] within deep learning networks can expose the subtle inconsistencies introduced by the inpainting operation in a different while complementary manner (refer to Fig. 1). Low-level features with high resolution have smaller receptive fields, thus primarily capturing signal artifacts, like edges and corners lacking semantic information [24]. Features at the middle level represent more abstract patterns and object parts, while high-level semantic features describing entire objects, scenes, or concepts are derived from the network’s deeper layers. Our goal is to fully leverage and enhance the diverse contributions of these features towards localizing image inpainting. To this end, we develop a new method for inpainting detection based on a Dense Feature Interaction Network (DeFI-Net). As shown in Table I, while some methods have recognized the importance of multi-level features for inpainting detection, simplistic interaction mechanisms struggle to capture complementary correlations across levels. DeFI-Net uses a feature pyramid architecture to capture and amplify multi-scale representations across various stages, which can more effectively reveal the feature-level interactions for detecting image inpainting. This structure adaptively directs the lower-level features, which carry edge and shape information, to refine the localization of manipulated edges while integrating the higher-level semantic features. Using DeFI-Net, we have developed a priori feature-guided edge attention module, incorporating inpainting edge prediction as an auxiliary task to enhance feature representation and improve localization accuracy at edge details. Evaluation on seven image inpainting datasets demonstrates the effectiveness of our approach, which achieves state-of-the-art performance in detecting inpainting across diverse models. The main contributions of our work can be summarized as follows:
@@
    We introduce the DeFI-Net to highlight how novel feature-level interactions aid inpainting detection, with a pyramid architecture hierarchically and thoroughly encoding multi-view features through dense connections.
@@
    We further design a complementary supervision approach by adding edge supervision with reverse attention on early-stage features to the inpainting region localization using multi-level features.
@@
    To effectively integrate mid-low-level details and high-level semantics features, we design a spatial weight learning module that adaptively assigns fusion weights to provide representative and compact contexts for recognizing inpainting.
@@
TABLE I: A comprehensive taxonomy of state-of-the-art methods for image inpainting detection.
  Methods Input information Backbone Multi-level feature Feature Interaction Mode
RGB Noise low-level mid-level high-level
MT-Net [25] + BayerConv2D, SRM filters VGG - - $\checkmark$ -
HP-FCN [17] - High-pass filters FCN - - $\checkmark$ -
IID-Net [18] + BayerConv2D, High-pass filters Customized cell (NAS) - - $\checkmark$ -
PSCC-Net [10] + - HRNet $\checkmark$ $\checkmark$ $\checkmark$ Progressive interaction
MVSS-Net [14] + BayerConv2D FCN $\checkmark$ $\checkmark$ $\checkmark$ Dominated by high-level features
PS-Net [20] + SRM filters, First-order derivative filters, Second-order derivative filters Dense-Net - - $\checkmark$ -
DiffForensics [15] + - U-Net $\checkmark$ $\checkmark$ $\checkmark$ Skip connections
CAT-Net [26] + DCT coefficient, Quantization table HRNet $\checkmark$ $\checkmark$ $\checkmark$ HRNet-based fusion
TruFor [27] + Noiseprint++ SegFormer $\checkmark$ $\checkmark$ $\checkmark$ Late concatenation
DeFI-Net(Ours) + - HRNet $\checkmark$ $\checkmark$ $\checkmark$ Dense interaction with context learning between layers
@@ II-A Image inpainting manipulation
Inpainting is a long-standing problem in image processing and computer vision [16]. The task is ill-posed because there is no unique solution based on the remaining part of the image, and various prior information is needed to hallucinate the missing region. Early inpainting methods using non-learning algorithms aimed to recover lost pixels using pixel redundancy and local continuity priors, prioritizing speed and simplicity. Based on the principle of local continuity, Bertalmio et al. [2] proposed a method that extends isobaths using fluid dynamics, achieving seamless mixing in complex structures. The Fast Marching Method [28] employs a level-set approach to efficiently propagate pixel values inward, enabling smooth transitions in small, regular inpainting regions. Herling and Broll [29] developed a mapping function to establish correspondence between missing and original regions of an image, incorporating a composite cost function to ensure spatial and appearance consistency. To enhance structural coherence, Huang et al. [30] integrated mid-level cues such as planes and regularities, making their approach particularly effective for structured scenes with architectural elements. Additionally, Guo et al. [31] applied Low-Rank Approximation to balance structural and textural recovery, utilizing low-rank decomposition to exploit image self-similarity and reconstruct both fine details and underlying structure. Collectively, these methods illustrate an evolution from basic pixel propagation techniques to advanced strategies incorporating structural integration, paving the way for more sophisticated inpainting tasks.
More recent effective inpainting methods are obtained with deep neural networks (DNNs) trained on large datasets [1], which allow for automatic feature learning and more context-aware inpainting results. Various types of DNN models have been explored, including the Generative Adversarial Network (GAN) [32, 33], Shift-Net [34] used U-Net architecture, recent Fast Fourier Convolution models [16, 35, 36, 37], and decoupled probabilistic modeling [38]. Learning-based methods emphasize adaptive feature selection and context integration to generate semantically coherent inpainting results. For example, Yu et al. [39] introduced gated convolutional layers to dynamically prioritize valid pixel information, enabling free-form inpainting across irregular mask shapes with high visual fidelity. Addressing challenges in large missing regions, where relying on neighboring pixel features alone is insufficient for accurate semantic reconstruction, Yu et al. [40] introduced the attention mechanism that dynamically learns global contextual information, resulting in semantically consistent outputs. To further enhance texture consistency, a two-stage inpainting strategy was proposed by Wu et al. [41], which combines a Local Binary Pattern network for structural estimation with spatial attention techniques. Additionally, to mitigate artificial artifacts and address mean and variance shift issues between missing and real areas, Yu et al. [42] introduced a region normalization process, which significantly improved inpainting quality.
@@ II-B Image inpainting detection
Image inpainting detection is often solved with machine learning based on convolutional neural networks (CNNs) [49]. These CNN-based approaches exploit hierarchical neural networks to automatically learn discriminative features for detecting inpainted regions. Zhu et al. [50] pioneered CNN-based image inpainting detection by employing a label matrix to guide automatic feature learning, and the weighted cross-entropy as the loss function. Several studies [51, 52, 25] employed Long Short-Term Memory (LSTM) models to learn artifacts for manipulation localization. HP-FCN [17] was proposed to enhance inpainting traces via high-pass filtering and learn discriminative features from image residuals. Similarly, IID-Net [18] used various combined filters to enhance inpainting features, and introduced a Neural Architecture Search (NAS) algorithm to adjust feature extraction cells. These conventional encoder-decoder structures without pooling layers may hinder the model’s capability to detect intricate tampering and result in spatial information loss. Several methods [27, 53, 11, 14] introduce multi-branch models to combine features from noise/frequency domains or patches [54]. New feature extraction models, including Transformer-based architectures [19, 55] or DenseNet-based approaches [56, 20] have been used for inpainting detection. Furthermore, researchers have explored multi-scale feature learning techniques. For example, MVSS-Net [13, 14] employed features learned from different stages of the backbone network and incorporated a two-branch supervision method considering RGB domain and noise patterns. Liu et al. [10] introduced PSCC-Net, leveraging the multi-stage features learned from the HRNet [21] for progressive mask prediction. GCA-Net [22] designed a gated context attention network to capture fine image discrepancies in inpainted images. To improve global context representations and minimize feature loss, they refer to the UNet++ [57] architecture, which utilizes multi-scale bottom-up feature propagation in the decoding process. Recently, Yu et al. introduced DiffForensics [15], which employs a self-supervised denoising diffusion paradigm integrated with an encoder-decoder architecture to detect and localize tampering.
@@
Figure 3: Architecture of the proposed DeFI-Net. Left: The DFPL module densely integrates multi-stage features extracted from the backbone network. Right: Structure of REAE and SA-FF modules for inpainting artifact enhancement and multi-level semantic fusion.
@@ III DeFI-Net
In this work, we introduce the DeFI-Net for image inpainting localization, as shown in Fig. 3. Taking an RGB image $I\in\mathbb{R}^{H\times W\times C}$ as input, the model will output a single-channel edge mask $M_{e}\in\mathbb{R}^{H\times W\times 1}$ and region mask $M_{r}\in\mathbb{R}^{H\times W\times 1}$. The model begins with multiple stages of feature extraction from the backbone network. The multi-stage features are interacted and fused by three components: 1) a Dense Feature Pyramid Learning structure (DFPL) to introduce multi-scale features containing detailed texture, contextual and semantic information (see Sec. III-A); 2) a Reverse Edge-Attention Enhancement (REAE), with mid-low-level features as a priori guiding the model to focus on edge inconsistencies caused by inpainting (see Sec. III-B); 3) a Spatial Adaptive Feature Fusion (SA-FF) aims to learn the fusion weights of the high-level and mid-low-level features adaptively, unifying the representation of inpainting to achieve accurate detection results ( see Sec. III-C).
@@ III-A Dense Feature Pyramid Learning
Different layers in a Convolutional Neural Network (CNN) model have receptive fields of different sizes and extract image features of different levels of detail. Generally, early or shadow layers mostly extract low-level and generic image cues (e.g., edges and corners), and the middle layers carry shapes and textural patterns. In contrast, deeper layers capture high-level semantics or specific image cues [60, 52, 24]. Artifacts of image inpainting may be reflected at all three levels of features, each contributing in a complementary manner. For example, pixel filling can cause distortions or blur at edges, detectable by low-level features. Middle-level features mainly capture texture inconsistencies in the filled areas, like structural mismatches or color discordance. Alternatively, high-level features identify changes in an image’s semantic content, such as missing object parts or altered scene interpretation, aiding the model in identifying inpainting areas.
Our goal is to design an effective extraction and fusion of these complementary features (as shown in Fig. 1) to identify image inpainting. We first use the backbone to extract multi-stage features from the input image. Using a CNN (e.g.,  [61]) divided into four stages as an example, we define the features from Stage 1 and Stage 2 as low-level features $f_{1}$ and $f_{2}$, the features extracted from Stage 3 as mid-level feature $f_{3}$, while the more abstract features from Stage 4 as high-level feature $f_{4}$. Inspired by the Feature Pyramid Network (FPN) [23] and Path Aggregation Network (PANet) [62], we design a multi-directional, multi-scale fusion method, as shown in Fig. 4. In addition, we treat multi-level features as independent entities, emphasizing the specificity of features at different stages. Our densely connected fusion approach improves the efficiency of information interaction between adjacent and cross layers, progressively enhancing context awareness.
Taking a single-scale feature as an example, our DFPL module considers the features of the same scale and all the neighboring features to fully exploit and enhance the useful context information for image inpainting localization. Specifically, to obtain the feature $f_{i}^{j}$ (where $i\in[1,4]$ denotes the stage of the backbone to which the current feature belongs, and $j\in[1,3]$ represents the state of the feature in DFPL, $j$ as 1 for the initial state involving channel reduction, 2 for the top-down path, and 3 for the bottom-up path), we concatenate its neighboring features in the channel dimension and use $1\times 1$ convolution to integrate and select features. This process can be expressed by the equation below, where we have removed the max pooling, upsampling operations and cases with $i$ in the boundary condition for simplicity.
Figure 4: Comparison of our DFPL with Feature Pyramid Network (FPN) [23] in multi-level feature interaction. Note that the displayed interaction process corresponds to the top-down path, as described in Eq. (1) with $j=2$.